The C-H stretching vibration in chloroform (CHCl3) occurs at 3000 cm-1
. Calculate the C–D stretching frequency in
CDCl3
1
Expert's answer
2020-04-27T01:51:36-0400

Let's assume that the C-H and C-D bonds can be regarded as isolated from the rest of the molecule. Then, the vibrational stretching frequency "\nu" is related to the reduced mass of the system of two atoms "\mu" :

"\nu = \frac{1}{2\pi}\sqrt{\frac{k}{\mu}}" ,

where "k" is the force constant of the chemical bond. Assuming that the force constant is the same for CHCl3 and CDCl3:

"\frac{\nu_H}{\nu_D} = \sqrt\frac{\mu_D}{\mu_H}"

The reduced masses of the systems are:

"\mu_H = \frac{m_Cm_H}{m_C+m_H} = \frac{12·1}{12 + 1} = 0.923" a.u.

"\mu_D = \frac{m_Cm_D}{m_C+m_D} = \frac{12·2}{12 + 2} = 1.71" a.u.

Therefore, the C-D stretching frequency is:

"\nu_D = 3000·\sqrt{\frac{\mu_H}{\mu_D}} = 2201 \text{ cm}^{-1}"

Answer: the C–D stretching frequency in CDCl3 is 2201 cm-1.